# Safety in Large Reasoning Models: A Survey

### v1

@@ 5.2 Inference-time Defenses for LRMs
Jiang et al. (2025) propose three decoding strategies, including ZeroThink, LessThink, and MoreThink, to verify model safety during reasoning. Making the reasoning safer at inference time could be a promising future direction, by verifying intermediate steps, filtering unsafe trajectories, or integrating reasoning-aware guard mechanisms during decoding.
@@ 5.3 Guard Models for LRMs
Through the proposed reasoning SFT and hard sample DPO, GuardReasoner (Liu et al., 2025b) is proposed to guide the guard model to deliberatively reason before making moderation decisions, improving performance, generalization ability, and explainability. Similarly, ThinkGuard (Wen et al., 2025) is developed via the proposed critique-augmented fine-tuning. X-Guard (Upadhayay et al., 2025) extends the reasoning-based guard model to the multi-lingual scenario.
@@ 7 Conclusion
This survey has comprehensively examined the emerging safety challenges posed by LRMs. We’ve identified unique vulnerabilities in these models that extend beyond traditional LLMs, mapping out the landscape of safety risks, adversarial attack vectors, and defense strategies. By organizing these elements into a detailed taxonomy, this work aims to facilitate future research that can enhance the security and reliability of these increasingly powerful AI systems while preserving their remarkable reasoning capabilities.

### v3

@@ 5.2 Inference-time Defenses for LRMs
Jiang et al. (2025) propose three decoding strategies, including ZeroThink, LessThink, and MoreThink, to verify model safety during reasoning. Making the reasoning safer at inference time could be a promising future direction, by verifying intermediate steps, filtering unsafe trajectories, or integrating reasoning-aware guard mechanisms during decoding. Wu et al. (2025a) introduce Thinking Intervention, a method that strategically injects guidance directly into the reasoning process to control model behavior and improve safety alignment without requiring additional training.
@@ 5.3 Guard Models for LRMs
Through the proposed reasoning SFT and hard sample DPO, GuardReasoner (Liu et al., 2025b) and GuardReasoner-VL (Liu et al., 2025d) are proposed to guide the guard model to deliberatively reason before making moderation decisions, improving performance, generalization ability, and explainability. Similarly, ThinkGuard (Wen et al., 2025) is developed via the proposed critique-augmented fine-tuning. X-Guard (Upadhayay et al., 2025) extends the reasoning-based guard model to the multi-lingual scenario.
@@ 7 Conclusion
This survey has comprehensively examined the emerging safety challenges posed by Large Reasoning Models. Through our analysis, we’ve identified several critical insights that distinguish LRM safety from traditional LLMs. First, LRMs expose their reasoning chains, creating new attack surfaces where adversaries can manipulate intermediate steps rather than just outputs, enabling sophisticated attacks like reasoning-based backdoors and hijacking that target the deliberative process itself. Second, traditional output-focused alignment methods prove insufficient for LRMs, as harmful reasoning can persist internally even when final outputs appear safe, necessitating novel approaches that consider the entire reasoning trajectory. These insights underscore the need for specialized safety research targeting LRMs, including standardized evaluation benchmarks for reasoning-specific vulnerabilities and human-in-the-loop alignment methods that can inspect and refine reasoning traces as these powerful models continue to advance into increasingly critical domains.